# From affine algebraic racks to Leibniz algebras

### v1

@@ 2.2. Pointed rack objects
2.2.1.
In the following, let $\pi_{1},\pi_{2}\colon Q\times Q\to Q$ be the projection morphisms, and let $\tau\colon Q\times Q\to Q\times Q$ denote the transposition $\tau=(\pi_{2},\pi_{1})$.
@@
Let $\mathcal{C}$ be a cartesian monoidal category with terminal object *. A pointed rack object or internal pointed rack in $\mathcal{C}$ is a quadruple $(Q,e,\triangleright,\triangleright^{-1})$ where
@@
  • $\triangleright$ and $\triangleright^{-1}$ are morphisms $Q\times Q\to Q$ called the rack operations of $Q$, and
@@
(2.1) $Q\times Q$$Q\times Q$$Q\times Q$$Q\times Q$$\scriptstyle{(\pi_{2},\,\triangleright^{-1})}$$\scriptstyle{(\triangleright,\pi_{1})}$$\scriptstyle{\operatorname{id}}$$\scriptstyle{(\triangleright,\pi_{1})}$$\scriptstyle{(\pi_{2},\,\triangleright^{-1})}$
(2.2) $Q\times Q\times Q$$Q\times Q\times Q$$Q\times Q$$Q{{Q\times Q}}$$\scriptstyle{(\triangleright,\pi_{1})\times\operatorname{id}}$$\scriptstyle{\mathord{\operatorname{id}}\times\triangleright}$$\scriptstyle{\mathord{\operatorname{id}}\times\triangleright}\scriptstyle{\triangleright}\scriptstyle{\triangleright}$  $Q\times Q\times Q$$Q\times Q\times Q$$Q\times Q$$Q{{Q\times Q}}$$\scriptstyle{\mathord{\operatorname{id}}\times(\pi_{2},\,\triangleright^{-1})}$$\scriptstyle{\triangleright^{-1}\times\operatorname{id}}$$\scriptstyle{\triangleright^{-1}\times\operatorname{id}}$$\scriptstyle{\triangleright^{-1}}$$\scriptstyle{\triangleright^{-1}}$
@@
  • idempotent or a quandle if $\triangleright\circ\Delta=\operatorname{id}=\triangleright^{-1}\circ\Delta$, where $\Delta\colon Q\to Q\times Q$ denotes the diagonal morphism $\Delta=(\operatorname{id},\operatorname{id})$.
@@
2.2.2.
As an important consequence of these definitions, we obtain automorphisms of pointed rack objects that generalize the left multiplication maps of pointed racks. Namely, let $Q$ be a pointed rack object in $\mathcal{C}$, and identify $\times Q\cong Q\cong Q\times*$. Given a global element of $Q$ (that is, a morphism $x\colon\ast\to Q$), we have automorphisms $L_{x},L_{x}^{-1}\colon Q\to Q$ given by
(2.4) $L_{x}\coloneq\triangleright\circ(x\times\operatorname{id}),\qquad L_{x}^{-1}\coloneq\triangleright^{-1}\circ(\mathord{\operatorname{id}}\times x).$
@@
Let $Q$ and $R$ be pointed rack objects in $\mathcal{C}$. We call $R$ a subrack of $Q$ if there exists a monomorphism of pointed rack objects $\iota\colon R\hookrightarrow Q$. We call $R$ a left ideal of $Q$ if there also exist morphisms $\widetilde{\triangleright}\colon Q\times R\to R$ and $\widetilde{\triangleright}^{-1}\colon R\times Q\to R$ such that the following diagrams commute:
@@
3. Algebraic racks and corack algebras
@@
3.1. Algebraic racks
Just as algebraic groups are defined to be group objects in $\mathsf{Sch}_{k}$}, we define algebraic racks to be pointed rack objects in $\mathsf{Sch}_{k}$}.
@@ 3.1. Algebraic racks
The category of algebraic racks is ${\mathsf{Rack}_{\ast}}({\mathsf{Sch}_{k}})$. Its objects and morphisms are called algebraic racks and morphisms of algebraic racks, respectively. The category of affine algebraic racks ${\mathsf{Rack}_{\ast}}({\mathsf{Aff}_{k}})$ is the full subcategory of ${\mathsf{Rack}_{\ast}}({\mathsf{Sch}_{k}})$ whose objects are affine.
More explicitly, an algebraic rack consists of a scheme $(Q,\mathcal{O}_{Q})$ and three morphisms of schemes $\triangleright,\triangleright^{-1}\colon Q\times Q\to Q$ and $e\colon\operatorname{Spec}(k)\to Q$ that make the diagrams in Definition 2.1 commute.
@@
3.1.1.
In the same way that we can think of algebraic groups as representable functors ${\mathsf{CAlg}_{k}}\to\mathsf{Set}$ that factor through $\mathsf{Grp}$, we can also think of algebraic racks as representable functors ${\mathsf{CAlg}_{k}}\to\mathsf{Set}$ that factor through $\mathsf{Rack}_{\ast}$}. This is the so-called functor of points approach that Grothendieck [grothendieck] introduced in 1973; see [milne]*Sec. 1a or any other modern algebraic geometry text for details.
@@
In particular, we call $\operatorname{OL}(\operatorname{GL}_{n},\mathbb{A}^{n})$ the omni-linear rack scheme and denote it by $\operatorname{OL}_{n}$. (See Proposition 7.1 for the motivation behind this name and notation.) As a functor, $\operatorname{OL}_{n}$ sends a $k$-algebra $R$ to the omni-linear rack $\operatorname{OL}(\operatorname{GL}_{n}(R),R^{n})$. Note that $\operatorname{OL}_{n}$ is an affine algebraic rack variety because $\operatorname{GL}_{n}$ and $\mathbb{A}^{n}$ are affine varieties. Indeed, $\operatorname{OL}_{n}$ shares the same underlying scheme but not the same morphisms as $\mathrm{Aff}_{n}$, the linear algebraic group of affine transformations of $\mathbb{A}^{n}$.
3.1.2.
Although the next example is uninteresting on its own, it obstructs adapting certain results about algebraic groups for algebraic racks.
@@
3.2. Corack algebras
Examples 3.3 and 3.4 suggest that in order to understand the coordinate ring of an affine algebraic rack $Q$, we have to study the pullbacks of the algebraic rack morphisms $\triangleright$, $\triangleright^{-1}$, and $e$. To that end, we introduce (commutative) corack algebras, which are to affine algebraic racks as Hopf algebras are to affine algebraic groups (see Proposition 4.1). Corack algebras are an important tool for studying affine algebraic racks and, indeed, proving Theorem 1.1.
3.2.1.
@@ 3.2. Corack algebras
3.2.2.
To make Definition 3.8 more concrete, let $A$ be a $k$-algebra with (commutative and associative) multiplication $\mu\colon A\otimes A\to A$, and let $\tau\colon A\otimes A\to A\otimes A$ denote the transposition $a\otimes b\mapsto b\otimes a$. A corack algebra structure on $A$ consists of two $k$-algebra homomorphisms $\nabla,\nabla^{-1}\colon A\to A\otimes A$ called the corack operations and a $k$-algebra homomorphism $\varepsilon\colon A\to k$ called the counit, coidentity, or augmentation satisfying the following five identities:
@@
    (Co-invertability) $(\mathord{\operatorname{id}}\otimes\mu)\circ(\tau\otimes\mu)\circ(\mathrm{\operatorname{id}}\otimes\nabla^{-1}\otimes\operatorname{id})\circ(\nabla\circ\operatorname{id})=(\mathord{\operatorname{id}}\otimes\operatorname{id})=(\mu\otimes\operatorname{id})\circ(\mu\otimes\tau)\circ(\mathord{\operatorname{id}}\otimes\nabla\otimes\operatorname{id})\circ(\mathord{\operatorname{id}}\otimes\nabla^{-1})$.
@@
    (Co-left distributivity) $(\mathord{\operatorname{id}}\otimes\nabla)\circ\nabla=(\mu\otimes\mathord{\operatorname{id}}\otimes\operatorname{id})\circ(\mathord{\operatorname{id}}\otimes\tau\otimes\operatorname{id})\circ(\nabla\otimes\nabla)\circ\nabla$.
@@
    (Co-right distributivity) $(\nabla^{-1}\otimes\operatorname{id})\circ\nabla^{-1}=(\mathord{\operatorname{id}}\otimes\mathord{\operatorname{id}}\otimes\mu)\circ(\mathord{\operatorname{id}}\otimes\tau\otimes\operatorname{id})\circ(\nabla^{-1}\otimes\nabla^{-1})\circ\nabla^{-1}$.
@@
    (Co-fixing) $(\varepsilon\otimes\operatorname{id})\circ\nabla=\operatorname{id}$.
@@
    (Co-fixedness) $(\mathord{\operatorname{id}}\otimes\varepsilon)\circ\nabla=\varepsilon$.
We identify $k\otimes A\cong A\cong A\otimes k$ in identities 4 and 5. Actually, this paper will only have occasion to use identities 2 and 4; the other three identities are less important for our purposes than the corresponding axioms of algebraic racks.
3.2.3.
Similarly, corack algebra homomorphisms are $k$-algebra homomorphisms that are compatible with $\nabla,\nabla^{-1}$, and $\varepsilon$ in the obvious ways. For example (cf. Section 4.1), if $\varphi\colon Q\to R$ is a morphism of affine algebraic racks, then the pullback $\varphi^{\sharp}\colon\mathcal{O}(R)\to\mathcal{O}(Q)$ satisfies
@@
Note that non-coassociativity means that the expression $(\nabla\otimes\operatorname{id})(\nabla f)=\sum f_{(1)(1)}\otimes f_{(1)(2)}\otimes f_{(2)}$ generally does not equal $(\mathord{\operatorname{id}}\otimes\nabla)(\nabla f)=\sum f_{(1)}\otimes f_{(2)(1)}\otimes f_{(2)(2)}$. This necessitates multi-indices, differing from the way one might use Sweedler notation for comultiplication in coalgebras.
Comment 3.10.
@@
As the name suggests, rack bialgebras are typically assumed to be coassociative but not associative, while corack algebras are associative but only coassociative in trivial cases (see Proposition 4.7). Also, rack bialgebras do not carry a second operation $\nabla^{-1}$.
@@ 4. Affine algebraic racks
For the remainder of the paper, we specialize our attention from algebraic racks ${\mathsf{Rack}_{\ast}}({\mathsf{Sch}_{k}})$ to affine algebraic racks ${\mathsf{Rack}_{\ast}}({\mathsf{Aff}_{k}})$.
@@ 4.1. Relationship with corack algebras
The relationship between affine algebraic racks and corack algebras mirrors the relationship between commutative Hopf algebras and affine algebraic groups (see, for example, [waterhouse]*Sec. 1.4). Namely, the morphisms $(\triangleright,\triangleright^{-1},e)$ of an affine algebraic rack $Q=\operatorname{Spec}(\mathcal{O}(Q))$ induce a canonical corack algebra structure $(\nabla,\nabla^{-1},\varepsilon)\coloneq(\triangleright^{\sharp},(\triangleright^{-1})^{\sharp},e^{\sharp})$ on the coordinate ring $\mathcal{O}(Q)$.
@@ 4.2. Worked examples
$\nabla=1\otimes\operatorname{id}.$
@@
$=\sum\Delta f_{(1)}(g,h)(Sf_{(2)})(g)$
$=\sum f_{(1)(1)}(g)f_{(1)(2)}(h)(Sf_{(2)})(g)$
$=\sum(f_{(1)(1)}Sf_{(2)})(g)f_{(1)(2)}(h).$
Hence,
$\nabla=(\mu\otimes\operatorname{id})\circ(\mathord{\operatorname{id}}\otimes S\otimes\operatorname{id})\circ(\mathord{\operatorname{id}}\otimes\tau)\circ(\Delta\otimes\operatorname{id})\circ\Delta,$
where $\mu\colon A\otimes A\to A$ denotes multiplication in $A$ and $\tau\colon A\otimes A\to A\otimes A$ denotes the transposition $a\otimes b\mapsto b\otimes a$.
@@
Let $n\geq 0$, and consider the omni-linear rack scheme $\operatorname{OL}_{n}$. As a scheme, we have $\operatorname{OL}_{n}=\operatorname{GL}_{n}\times\mathbb{A}^{n}$, so the underlying $k$-algebra of the corack algebra $\mathcal{O}(\operatorname{OL}_{n})$ is

### v2

@@ 3. Pointed rack objects
3.2.
In the following, recall that $\pi_{1},\pi_{2}\colon Q^{2}\to Q$ denote the projection morphisms, and recall that $\tau\colon Q^{2}\xrightarrow{\sim}Q^{2}$ denotes the transposition $\tau=(\pi_{2},\pi_{1})$.
@@
Let $\mathcal{C}$ be a cartesian monoidal category with terminal object *. A pointed rack object or internal pointed rack in $\mathcal{C}$ is a quadruple $(Q,e,\triangleright,\triangleright^{-1})$ (sometimes denoted only by $(Q,\triangleright)$ or $Q$) where
@@
  • $\triangleright$ and $\triangleright^{-1}$ are morphisms $Q^{2}\to Q$ called the rack operations of $Q$, and
@@
(3.1) $Q^{2}$$Q^{2}$$Q^{2}$$Q^{2}$$\scriptstyle{(\pi_{2},\,\triangleright^{-1})}$$\scriptstyle{(\triangleright,\pi_{1})}$$\scriptstyle{\mathrm{id}}$$\scriptstyle{(\triangleright,\pi_{1})}$$\scriptstyle{(\pi_{2},\,\triangleright^{-1})}$
[…]
@@
  • idempotent or a quandle if $\triangleright\circ\Delta=\mathrm{id}=\triangleright^{-1}\circ\Delta$, where $\Delta\colon Q\to Q^{2}$ denotes the diagonal morphism $\Delta=(\mathrm{id},\mathrm{id})$.
@@
Note that every pointed rack also satisfies diagrams similar to those in (3.3) that generalize the inverse fixing and inverse fixedness identities 4 and 5.
@@
$3.$
As in [rack-roll], we can also define rack objects and morphisms of rack objects in cartesian monoidal categories $\mathcal{C}$ by omitting all axioms relating to units $e$ in Definitions 3.1 and 3.2. Let $\mathsf{Rack}(\mathcal{C})$ denote the category of rack objects in $\mathcal{C}$.
Similarly to Kinyon [kinyon], we are mainly interested in pointed rack objects ${\mathsf{Rack}_{\ast}}({\mathsf{Aff}_{k}})$ in the category of affine schemes instead of rack objects $\mathsf{Rack}({\mathsf{Aff}_{k}})$. This is because the unit $e$ gives us a canonical choice of tangent space $\mathfrak{q}\coloneq T_{e}Q$ on which to define a Leibniz algebra structure.
3.4.
As an important consequence of Definitions 3.1 and 3.2, we obtain automorphisms of pointed rack objects that generalize the left multiplication maps of pointed racks. Namely, let $Q$ be a pointed rack object in $\mathcal{C}$, and identify $\times Q\cong Q\cong Q\times*$. Given a global element of $Q$ (that is, a morphism $x\colon\ast\to Q$), we have automorphisms $L_{x},L_{x}^{-1}\colon Q\to Q$ given by
(3.4) $L_{x}\coloneq\triangleright\circ(x\times\mathrm{id}),\qquad L_{x}^{-1}\coloneq\triangleright^{-1}\circ(\mathord{\mathrm{id}}\times x).$
@@
Let $Q$ and $R$ be pointed rack objects in $\mathcal{C}$. We call $R$ a subrack of $Q$ if there exists a monomorphism of pointed rack objects $\iota\colon R\hookrightarrow Q$. We call $R$ a left ideal of $Q$ if there also exist morphisms $\widetilde{\triangleright}\colon Q\times R\to R$ and $\widetilde{\triangleright}^{-1}\colon R\times Q\to R$ such that the following diagrams two commute:
@@
4. Algebraic racks
@@
4.1.
Just as algebraic groups are defined to be group objects in $\mathsf{Sch}_{k}$}, we define algebraic racks to be pointed rack objects in $\mathsf{Sch}_{k}$}. For the purposes of proving Theorem 1.1, we are primarily interested in working over a ground field $k$, but the definitions work when $k$ is replaced with an arbitrary commutative ring $S$.
@@
The categories of algebraic racks and affine algebraic racks are ${\mathsf{Rack}_{\ast}}({\mathsf{Sch}_{k}})$ and ${\mathsf{Rack}_{\ast}}({\mathsf{Aff}_{k}})$, respectively.
Definition 4.2.
The categories of rack schemes and affine rack schemes are $\mathsf{Rack}({\mathsf{Sch}_{k}})$ and $\mathsf{Rack}({\mathsf{Aff}_{k}})$, respectively.
More explicitly, an algebraic rack consists of a scheme $(Q,\mathcal{O}_{Q})$ and three morphisms of schemes $\triangleright,\triangleright^{-1}\colon Q^{2}\to Q$ and $e\colon\operatorname{Spec}(k)\to Q$ that make the diagrams in Definition 3.1 commute.
@@
4.2.
In the same way that we can consider every (affine) algebraic group as a (representable) functor ${\mathsf{CAlg}_{k}}\to\mathsf{Set}$ that factors through $\mathsf{Grp}$, we can also consider every (affine) algebraic rack as a (representable) functor ${\mathsf{CAlg}_{k}}\to\mathsf{Set}$ that factors through $\mathsf{Rack}_{\ast}$}. This is the so-called functor of points approach that Grothendieck [grothendieck] introduced in 1973; see [milne]*Sec. 1a or any other modern algebraic geometry text for details.
@@
In particular, we call $\operatorname{OL}_{n}\coloneq\operatorname{OL}(\operatorname{GL}_{n},\mathbb{A}^{n})$ the omni-linear rack variety. (See Proposition 13.1 for the motivation behind this name and notation.) As a functor, $\operatorname{OL}_{n}$ sends every $k$-algebra $R$ to the omni-linear rack $\operatorname{OL}(\operatorname{GL}_{n}(R),R^{n})$. Note that $\operatorname{OL}_{n}$ is an affine algebraic rack variety because $\operatorname{GL}_{n}$ and $\mathbb{A}^{n}$ are affine varieties. Indeed, $\operatorname{OL}_{n}$ shares the same underlying scheme but not the same morphisms as $\mathrm{Aff}_{n}$, the linear algebraic group of affine transformations of $\mathbb{A}^{n}$.
4.3.
Although the next example is uninteresting on its own, it is an important obstruction to adapting certain results about algebraic groups for algebraic racks.
@@
If we turn our attention to rack schemes $\mathsf{Rack}({\mathsf{Sch}_{k}})$ rather than algebraic racks ${\mathsf{Rack}_{\ast}}({\mathsf{Sch}_{k}})$, then Example 4.6 generalizes to an even broader class of rack schemes.
Example 4.8.
Let $Q$ be any scheme, and let $\varphi$ be an automorphism of $Q$. Define morphisms $\triangleright,\triangleright^{-1}\colon Q^{2}\to Q$ by
$\triangleright\coloneq\varphi\circ\pi_{2},\qquad\triangleright^{-1}\coloneq\varphi^{-1}\circ\pi_{1},$
where $\pi_{1},\pi_{2}\colon Q^{2}\to Q$ are the projection morphisms. Then $Q$ is a rack scheme called a permutation rack scheme. If $Q$ is nonempty and $\varphi=\mathrm{id}_{Q}$, then we recover Example 4.6.
5. Corack algebras
Examples 4.4 and 4.5 suggest that in order to understand the coordinate algebra $A$ of an affine algebraic rack $Q=\operatorname{Spec}(A)$, we have to study the pullbacks of the algebraic rack morphisms $\triangleright$, $\triangleright^{-1}$, and $e$. To that end, this section introduces (commutative) corack algebras, which are to affine algebraic racks as commutative Hopf algebras are to affine algebraic groups (see Proposition 6.1). Accordingly, corack algebras allow us to both state and prove Theorem 1.1.
5.1.
@@ 5. Corack algebras
5.2.
To make Definition 5.2 more concrete, let $A$ be a $k$-algebra with (commutative and associative) multiplication $\mu\colon A\otimes A\to A$, and recall that $\tau$ denotes the transposition $a\otimes b\mapsto b\otimes a$.
A corack algebra structure on $A$ consists of two $k$-algebra homomorphisms $\nabla,\nabla^{-1}\colon A\to A\otimes A$ called the corack operations and a $k$-algebra homomorphism $\varepsilon\colon A\to k$ called the counit, coidentity, or augmentation that satisfy the following five identities induced by the diagrams in (3.1)–(3.3):
@@
    (Co-invertability) $(\mathord{\mathrm{id}}\otimes\mu)\circ(\tau\otimes\mu)\circ(\mathrm{\mathrm{id}}\otimes\nabla^{-1}\otimes\mathrm{id})\circ(\nabla\circ\mathrm{id})=(\mathord{\mathrm{id}}\otimes\mathrm{id})=(\mu\otimes\mathrm{id})\circ(\mu\otimes\tau)\circ(\mathord{\mathrm{id}}\otimes\nabla\otimes\mathrm{id})\circ(\mathord{\mathrm{id}}\otimes\nabla^{-1})$.
@@
    (Co-left distributivity) $(\mathord{\mathrm{id}}\otimes\nabla)\circ\nabla=(\mu\otimes\mathord{\mathrm{id}}\otimes\mathrm{id})\circ(\mathord{\mathrm{id}}\otimes\tau\otimes\mathrm{id})\circ(\nabla\otimes\nabla)\circ\nabla$.
@@
    (Co-right distributivity) $(\nabla^{-1}\otimes\mathrm{id})\circ\nabla^{-1}=(\mathord{\mathrm{id}}\otimes\mathord{\mathrm{id}}\otimes\mu)\circ(\mathord{\mathrm{id}}\otimes\tau\otimes\mathrm{id})\circ(\nabla^{-1}\otimes\nabla^{-1})\circ\nabla^{-1}$.
@@
    (Co-fixing) $(\varepsilon\otimes\mathrm{id})\circ\nabla=\mathrm{id}$.
@@
    (Co-fixedness) $(\mathord{\mathrm{id}}\otimes\varepsilon)\circ\nabla=\varepsilon$.
We also obtain the following two identities:
  1. (C4’)
    (Co-inverse fixing) $(\mathrm{id}\otimes\varepsilon)\circ\nabla^{-1}=\mathrm{id}$.
  2. (C5’)
    (Co-inverse fixedness) $(\varepsilon\otimes\mathrm{id})\circ\nabla^{-1}=\varepsilon$.
We identify $k\otimes A\cong A\cong A\otimes k$ in identities 4, 5, 4, and 5. We identify $k$ as a $k$-subalgebra of $A$ in identities 5 and 5.
5.3.
Similarly, corack algebra homomorphisms are $k$-algebra homomorphisms that are compatible with the structure morphisms $\nabla,\nabla^{-1}$, and $\varepsilon$ in the obvious ways. For example (cf. Section 6.1), if $\varphi\colon Q\to R$ is a morphism of affine algebraic racks, then the pullback $\varphi^{\sharp}\colon\mathcal{O}(R)\to\mathcal{O}(Q)$ satisfies
@@
Note that non-coassociativity means that the expression $(\nabla\otimes\mathrm{id})(\nabla f)=\sum f_{(1)(1)}\otimes f_{(1)(2)}\otimes f_{(2)}$ generally does not equal $(\mathord{\mathrm{id}}\otimes\nabla)(\nabla f)=\sum f_{(1)}\otimes f_{(2)(1)}\otimes f_{(2)(2)}$. This necessitates multi-indices, differing from the way one might use Sweedler notation for comultiplication in coalgebras.
Comment 5.4.
@@
As the name suggests, rack bialgebras are typically assumed to be coassociative but not associative, while corack algebras are associative but only coassociative in trivial cases (see Proposition 6.8). Also, rack bialgebras are not assumed to carry a second operation $\nabla^{-1}$.
@@ 6. Affine algebraic racks
In the next four sections, we specialize our attention from algebraic racks ${\mathsf{Rack}_{\ast}}({\mathsf{Sch}_{k}})$ to affine algebraic racks ${\mathsf{Rack}_{\ast}}({\mathsf{Aff}_{k}})$.
@@ 6.1. Relationship with corack algebras
The relationship between affine algebraic racks and corack algebras mirrors the relationship between commutative Hopf algebras and affine algebraic groups (see, for example, [waterhouse]*Sec. 1.4). Namely, the morphisms $(\triangleright,\triangleright^{-1},e)$ of an affine algebraic rack $Q=\operatorname{Spec}(\mathcal{O}(Q))$ induce a canonical corack algebra structure $(\nabla,\nabla^{-1},\varepsilon)\coloneq(\triangleright^{\sharp},(\triangleright^{-1})^{\sharp},e^{\sharp})$ on the coordinate algebra $\mathcal{O}(Q)$.
@@ 6.2. Worked examples
$\nabla=1\otimes\mathrm{id}.$
Remark 6.5.
If $(Q,\triangleright)$ is an affine rack scheme (as opposed to an affine algebraic rack, which is equipped with a unit $e$), then we can still define $(\nabla,\nabla^{-1})\coloneq(\triangleright^{\sharp},(\triangleright^{-1})^{\sharp})$.
In particular, if $(Q,\triangleright)$ is a permutation rack scheme with respect to $\varphi\in\operatorname{Aut}(Q)$ as in Example 4.8, then the same calculation as in Example 6.4 yields
$\nabla=1\otimes\varphi^{\sharp}.$
@@
$=\sum\Delta f_{(1)}(g,h)(S(f_{(2)}))(g)$
$=\sum f_{(1)(1)}(g)f_{(1)(2)}(h)(S(f_{(2)}))(g)$
$=\sum(f_{(1)(1)}S(f_{(2)}))(g)f_{(1)(2)}(h).$
Since $\Delta$ is coassociative, it follows that
$\nabla f=\sum f_{(1)}S(f_{(3)})\otimes f_{(2)}.$
Equivalently, let $\mu\colon A\otimes A\to A$ denote multiplication in $A$, and let $\tau$ denote the transposition $a\otimes b\mapsto b\otimes a$. Then
$\nabla=(\mu\otimes\mathrm{id})\circ(\mathord{\mathrm{id}}\otimes S\otimes\mathrm{id})\circ(\mathord{\mathrm{id}}\otimes\tau)\circ(\Delta\otimes\mathrm{id})\circ\Delta.$
@@
Let $n\geq 0$, and consider the omni-linear rack variety $\operatorname{OL}_{n}$. As an affine variety, we have $\operatorname{OL}_{n}=\operatorname{GL}_{n}\times\mathbb{A}^{n}$, so the underlying $k$-algebra of the corack algebra $\mathcal{O}(\operatorname{OL}_{n})$ is